OpenStudy (anonymous):

If the dimensions of a rectangle become one-fifth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle. Enter your answer in numerical form, using the / for the fraction bar if needed.

OpenStudy (anonymous):

please. i need help.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

these people should help you

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

okay, thank you

OpenStudy (anonymous):

fan those people to

OpenStudy (amistre64):

area of a rectangle: bh area of 1/5th = 1/5 b 1/5 h = (1/5)^2 bh

OpenStudy (akashdeepdeb):

Okay so in this question let us say we have Length l and breadth b So the area = l times b Now if it becomes 1/5 Then length = l/5 breadth = b/5 So the new area will be l/5 times b/5 that is lb/25 So the new area would be 1/25 of the original area!! Got it? :)

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

and yes i got it

OpenStudy (akashdeepdeb):

:)